please help me solve this expression 5x-8x
please help me solve this expression -3a+a
please help me solve this expressions 5x+8x+3x
please help me solve this expression 6a-4a+5a

Answer by amine(11)

5x-8x=-3x
-3a+a=-2a
5x+8x+3x=16x
6a-4a+5a=7a